The 614𝐶 isotope has a half-life of 5730 yr. If at some time a sample contains 1.00 x 1022 carbon-14 nuclei, what is the activity of the sample? 


1
Expert's answer
2021-05-18T11:09:59-0400

Given

Number of nuclei is "N= 1.00\times 10^{22}"

Half life is "T_{1/2}=5730 year=1.807 \times 10^{11} seconds"

The activity of the sample is

"R=\lambda N
 =\frac{ln2}{t_{1/2}}N
 =\frac{ln2}{1.807 \times 10^{11} seconds}( 1.00\times 10^{22} )"

"=3.836\times 10^{10} decay/s"